Is there a math formula that relates Doppler effect and speed of the moving source? And is it different when applied to sound or when applied to light?
1
Expert's answer
2019-09-04T09:31:24-0400

For sound:

"\omega=\omega_0\frac{1+\frac{u}{c}}{1-\frac{v}{c}}"

u — is the receiver velocity relative to the medium.

v — is the wave propagation velocity in the medium.

c — is the wave propagation velocity in the medium.

w0 —— is the source wave frequency.

For light:

"\omega=\omega_0\frac{\sqrt{1-\frac{v^2}{c^2}}}{1+\frac{v}{c}cos\theta}"

c - is the velocity of light.

v - is the source speed relative to receiver.

"\theta" - is the angle between the direction of the source and the velocity vector associated with the reference frame of the receiver.

w - is the frequency that the receiver will fix.